# Light and heavy $\Lambda$ hyperclusters in nuclear matter with RMF models

### v1

@@ III.1 $\Lambda$ hypernuclei
To show these more explicitly, in Fig. 2 we present the obtained density profiles of 4He, ${}^{5}_{\Lambda}$e, and ${}^{6}_{\Lambda\Lambda}$e in vacuum with $n_{\mathrm{gas}}=0$ and the Coulomb potential $A=eN_{p}/4\pi r$ outside of the nucleus. The density profile of $\Lambda$ hyperons for ${}^{6}_{\Lambda\Lambda}$e is about twice of ${}^{5}_{\Lambda}$e since they both occupy the same orbital $1s_{1/2}$, while the density of each $\Lambda$ hyperon at the center becomes slightly larger for ${}^{6}_{\Lambda\Lambda}$e. Due to the presence of $\Lambda$ hyperons, the core 4He becomes more compact with its central density increased, where $n_{p}(0)+n_{n}(0)=0.161$, 0.182, and 0.204 fm-3 for 4He, ${}^{5}_{\Lambda}$e, and ${}^{6}_{\Lambda\Lambda}$e. This destabilizes the core and leads to a reduction of $B_{\Lambda}\approx 2.7$ MeV by approximately 1 MeV from that of the single-particle energy for $\Lambda$ hyperon ($\varepsilon_{1s_{1/2}}\approx-4$ MeV) in ${}^{5}_{\Lambda}$e. The corresponding $\Lambda$-$\Lambda$ interaction energy for ${}^{6}_{\Lambda\Lambda}$e is $\Delta B_{\Lambda\Lambda}=-0.133$ MeV and indicates a repulsive $\Lambda$-$\Lambda$ interaction, while in fact the $\Lambda$-$\Lambda$ interaction is attractive with $\Delta B_{\Lambda\Lambda}=0.67\pm 0.17$ MeV according to the KEK hybrid-emulsion experiment E373 Takahashi et al. (2001); Ahn et al. (2013). This can be fixed by introducing explicitly $\Lambda$-$\Lambda$ attractive interaction via $\sigma^{*}$ meson exchange, which is nonetheless neglected here since $\Delta B_{\Lambda\Lambda}$ is relatively small compared with $B_{\Lambda\Lambda}$ ($=6.91\pm 0.16$ MeV) and we only consider systems with $N_{\Lambda}\leq 2$.
@@ III.2 $\Lambda$ hyperclusters in nuclear medium
We then investigate the properties of various (hyper)clusters in nuclear medium. In Fig. 3 the binding energy of H and He clusters emersed in nuclear medium with density $n_{\mathrm{gas}}=n_{p,\mathrm{gas}}+n_{n,\mathrm{gas}}$ and proton fraction $Y_{p}=n_{p,\mathrm{gas}}/n_{\mathrm{gas}}$ are presented, which are obtained by employing the relativistic density functional DD-LZ1 Wei et al. (2020). The proton fraction $Y_{p}$ is thus related to the asymmetry parameter of the nuclear matter by $\delta=(n_{n,\mathrm{gas}}-n_{p,\mathrm{gas}})/n_{\mathrm{gas}}=1-2Y_{p}$. At vanishing densities $n_{\mathrm{gas}}\rightarrow 0$, as indicated by $B_{0}$ in Table 3, in-medium clusters become more bound than those in vacuum due to the reduction of Coulomb energy Xia (2025). The inclusion of $\Lambda$ hyperons also increase the binding energy by $B_{\Lambda}$ and $B_{\Lambda\Lambda}$ for single and double $\Lambda$ hyperclusters. As $n_{\mathrm{gas}}$ increases, the binding energies of clusters decrease, which would eventually become unstable and melt at sufficient large densities, i.e., Mott transition. Despite the increased binding energy for $\Lambda$ hyperclusters at small $n_{\mathrm{gas}}$, the critical densities for vanishing binding energy $B=0$ are smaller for $\Lambda$ hyperclusters than that of normal clusters without $\Lambda$ hyperons. This is attributed to the reduction of binding energy per baryon for light nuclei with addition of $\Lambda$ hyperons, i.e., $B_{\mathrm{core}}/(A-1)>B_{\Lambda}$ as indicated in Fig. 1. Meanwhile, the variation of binding energies with respect to $Y_{p}$ are found to be sensitive to the differences between neutron and proton numbers in the cluster. For (hyper)clusters 3H, ${}^{4}_{\Lambda}$, and ${}^{5}_{\Lambda\Lambda}$, their binding energies decrease with the proton fraction of nuclear medium $Y_{p}$. If we replace a neutron in 3H by a $\Lambda$ hyperon, as indicated in Fig. 3, the binding energy for ${}^{3}_{\Lambda}$ is generally insensitive to $Y_{p}$ since $N_{p}=N_{n}=1$. Similarly, the binding energy for 4He, ${}^{5}_{\Lambda}$e, and ${}^{6}_{\Lambda\Lambda}$e also vary little with respect to the variation of $Y_{p}$ since $N_{p}=N_{n}=2$.
@@
${}^{3}_{\Lambda}$ 5.208 --3.380 $195$-0.0131 --0.31 $0.49$
3H 11.53 --3.089 $113$-1.11 --1.2 $1.1$
${}^{4}_{\Lambda}$ 13.48 --3.822 $144$-1.19 --1.1 $1.2$
${}^{5}_{\Lambda\Lambda}$ 15.31 --4.580 $170$-1.19 --1.1 $1.2$
3He 10.57 --4.388 $149$ $1.197$-0.26 $1.0$
4He 27.70 --5.029 $122$ $0.166$-0.60 $0.77$
${}^{5}_{\Lambda}$e 30.39 --5.800 $141$ $0.233$-0.53 $0.51$
${}^{6}_{\Lambda\Lambda}$e 32.97 --6.577 $162$ $0.285$-0.47 $0.30$
8Be 43.34 --10.06 $262$ $0.268$ $100$
${}^{9}_{\Lambda}$e 50.80 --10.66 $230$ $0.438$ $0490$
12C 95.40 --14.55 $221$ $1.03$-0.37 0
${}^{13}_{\Lambda}$ 106.9 --15.44 $249$ $1.03$-0.34 0
16O 134.6 --19.32 $337$ $1.41$-0.44 0
${}^{17}_{\Lambda}$ 145.6 --20.19 $337$ $1.47$-0.47 0
40Ca 383.0 --48.47 $714$ $4.53$-1.3 0
${}^{41}_{\Lambda}$a 399.7 --49.27 $723$ $4.69$-1.5 0
48Ca 454.2 --52.32 $712$-4.81 --3.4 0
${}^{49}_{\Lambda}$a 473.1 --53.08 $715$-4.61 --3.5 0
@@
The best-fit coefficients are then indicated in Table 3. The coefficient $B_{0}$ represent the binding energy of (hyper)clusters in nuclear medium at vanishing $n_{\mathrm{gas}}$, which is slightly more bound than those in vacuum due to the reduction of Coulomb energy. The coefficients $a$ and $b$ indicate the density dependent behavior for the binding energy of clusters in nuclear medium. Despite the increased binding energy with the inclusion of $\Lambda$ hyperons, $a$ becomes larger and leads to a faster decline for the binding energy as $n_{\mathrm{gas}}$ increases. The coefficients $c$, $d$, and $f$ are connected to the variation of binding energy with respect to the isospin asymmetry of nuclear medium, which are dominated by the differences between neutron and proton numbers in the cluster, e.g., the binding energies for ${}^{3}_{\Lambda}$, 4He, ${}^{5}_{\Lambda}$e, and ${}^{6}_{\Lambda\Lambda}$e are roughly insensitive to $Y_{p}$ while those of 3H, ${}^{4}_{\Lambda}$, and ${}^{5}_{\Lambda\Lambda}$ decrease with $Y_{p}$ as $c$ takes positive values. This is mainly attributed to the isovector effects in nuclear medium with the differences in the mean-field potentials of protons and neutrons change as we vary $Y_{p}$.
Figure 4: Binding energy per nucleon for various (hyper)clusters emersed in SYM and PNM as functions of the density of nuclear medium $n_{\mathrm{gas}}$, which are obtained by employing the relativistic density functional DD-LZ1 Wei et al. (2020). The thin curves indicate normal clusters and thick curves correspond to those with additional one $\Lambda$ hyperon.
In Fig. 4 we present the binding energy per baryon for heavier (hyper)clusters emersed in SYM and pure neutron matter (PNM, $Y_{p}=0$) as functions of $n_{\mathrm{gas}}$. Similar to light clusters, the binding energies of (hyper)clusters decrease with $n_{\mathrm{gas}}$. Nevertheless, since the binding energy per baryon increase for $\Lambda$ hyperclusters with $N_{p}\geq 4$, they become more bound with the Mott transition densities increased. By comparing the upper and lower panels in Fig. 4, the variations of binding energies with respect to $Y_{p}$ can be identified, where ($\Lambda$ hyper)clusters with $N_{p}=N_{n}\lesssim 6$ are generally insensitive to the variations in $Y_{p}$. By fitting the binding energies indicated in Fig. 4 with Eq. (36), we can fix the corresponding coefficients, which are indicated in Table 3. The variations of binding energies with respect to $n_{\mathrm{gas}}$ and $Y_{p}$ can then be reproduced.
@@
The charge radii $R_{\mathrm{ch}}$ for the (hyper)clusters as functions of the density of nuclear medium $n_{\mathrm{gas}}$ are presented in Fig. 5, which correspond to those in Fig. 4 and are obtained with the relativistic density functional DD-LZ1 Wei et al. (2020). It is found that $R_{\mathrm{ch}}$ are generally increasing with $n_{\mathrm{gas}}$ while the binding energy is decreasing. Similar to the cases in vacuum as indicated in Table 2, the addition of a $\Lambda$ hyperon reduces the charge radii. The slope of $R_{\mathrm{ch}}$ as a function of $n_{\mathrm{gas}}$ also becomes smaller with the addition of a $\Lambda$ hyperon. By comparing the upper and lower panels in Fig. 5, the isovector effects in nuclear medium can be identified, e.g., $R_{\mathrm{ch}}$ increases slightly faster for (hyper)clusters with in PNM than those in SYM as we increase $n_{\mathrm{gas}}$. For heavy (hyper)clusters such as 48Ca and ${}^{48}_{\Lambda}$a, the variations of $R_{\mathrm{ch}}$ with respect to $Y_{p}$ are less significant.
@@
Figure 6: Binding energies of 3H (thin curves), 4He (thin curves), ${}^{4}_{\Lambda}$ (thick curves) and ${}^{5}_{\Lambda}$e (thick curves) emersed in SYM and PNM as functions of $n_{\mathrm{gas}}$, which are obtained adopting the relativistic density functionals PK1 Long et al. (2004), TM1 Sugahara and Toki (1994b), and DD-LZ1 Wei et al. (2020). The results are fitted with Eq. (36) with the coefficients indicated in Table 4.
@@
Finally, to show the variations due to the uncertainties in relativistic density functionals, we examine the binding energies of 3H, 4He, ${}^{4}_{\Lambda}$, and ${}^{5}_{\Lambda}$e emersed in SYM and PNM adopting the functionals PK1 Long et al. (2004), TM1 Sugahara and Toki (1994b), and DD-LZ1 Wei et al. (2020), which are presented in Fig. 6. As expected, the binding energy increases with the inclusion of $\Lambda$ hyperon, while the binding energy per baryon decreases since $B_{\mathrm{core}}/(A-1)>B_{\Lambda}$ for light hyperclusters as indicated in Fig. 1. The binding energies of the (hyper)clusters are generally decreasing with the density of nuclear matter $n_{\mathrm{gas}}$. For the cluster 4He, the binding energy decline slowly with $n_{\mathrm{gas}}$, while slight variations on the in-medium binding energy shift are observed when we adopt different functionals. Due to the non-linear self-couplings of mesons, the binding energies of 3H, ${}^{4}_{\Lambda}$, and ${}^{5}_{\Lambda}$e decrease drastically at $n_{\mathrm{gas}}\approx 0.0005$-0.0011 $\mathrm{fm}^{-3}$ when we adopt non-linear relativistic density functionals PK1 Long et al. (2004) and TM1 Sugahara and Toki (1994b), while the density-dependent functional DD-LZ1 Wei et al. (2020) predicts relatively slow decline for the binding energies. It is worth mentioning that the binding energy of ${}^{5}_{\Lambda}$e decreases drastically when non-linear relativistic density functionals are adopted, even though that of 4He decreases slowly with $n_{\mathrm{gas}}$. In such cases, the addition of $\Lambda$ hyperons destabilize the light (hyper)clusters which is particular the case when non-linear relativistic density functionals are adopted.
@@
Table 4: Coefficients for the expansion formula (36) of binding energies for in-medium 3H, 4He, ${}^{4}_{\Lambda}$, and ${}^{5}_{\Lambda}$e presented in Fig. 6, where $f=0$ is assumed.

### v2

@@ III.1 $\Lambda$ hypernuclei
To show these more explicitly, in Fig. 2 we present the obtained density profiles of 4He, $^{5}_{\Lambda}$He, and $^{6}_{\Lambda\Lambda}$He in vacuum with $n_{\mathrm{gas}}=0$ and the Coulomb potential $A=eN_{p}/4\pi r$ outside of the nucleus. The density profile of $\Lambda$ hyperons for $^{6}_{\Lambda\Lambda}$He is about twice higher than that of $^{5}_{\Lambda}$He since they both occupy the same orbital $1s_{1/2}$, while the density of each $\Lambda$ hyperon at the center becomes slightly larger for $^{6}_{\Lambda\Lambda}$He. Due to the presence of $\Lambda$ hyperons, the core 4He becomes more compact with its increasing central density, where $n_{p}(0)+n_{n}(0)=0.161$, 0.182, and 0.204 fm-3 for 4He, $^{5}_{\Lambda}$He, and $^{6}_{\Lambda\Lambda}$He. This destabilizes the core and leads to a reduction of $B_{\Lambda}\approx 2.7$ MeV by approximately 1 MeV from that of the single-particle energy for $\Lambda$ hyperon ($\varepsilon_{1s_{1/2}}\approx-4$ MeV) in $^{5}_{\Lambda}$He. The corresponding $\Lambda$-$\Lambda$ interaction energy for $^{6}_{\Lambda\Lambda}$He is $\Delta B_{\Lambda\Lambda}=-0.133$ MeV and indicates a repulsive $\Lambda$-$\Lambda$ interaction, while in fact the $\Lambda$-$\Lambda$ interaction is attractive with $\Delta B_{\Lambda\Lambda}=0.67\pm 0.17$ MeV according to the KEK hybrid-emulsion experiment E373 Takahashi et al. (2001); Ahn et al. (2013). This can be fixed by introducing explicitly $\Lambda$-$\Lambda$ attractive interaction via $\sigma^{*}$ meson exchange Rong et al. (2021), which is nonetheless neglected here since $\Delta B_{\Lambda\Lambda}$ is relatively small compared with $B_{\Lambda\Lambda}$ ($6.91\pm 0.16$ MeV) and we only consider systems with $N_{\Lambda}\leq 2$.
@@ III.2 $\Lambda$ hyperclusters in nuclear medium
We then investigate the properties of various (hyper)clusters in nuclear medium. In Fig. 3 the binding energy of H and He clusters emersed in nuclear medium with density $n_{\mathrm{gas}}=n_{p,\mathrm{gas}}+n_{n,\mathrm{gas}}$ and proton fraction $Y_{p}=n_{p,\mathrm{gas}}/n_{\mathrm{gas}}$ are presented, which are obtained by employing the relativistic density functional DD-LZ1 Wei et al. (2020). The proton fraction $Y_{p}$ is thus related to the asymmetry parameter of the nuclear matter by $\delta=(n_{n,\mathrm{gas}}-n_{p,\mathrm{gas}})/n_{\mathrm{gas}}=1-2Y_{p}$. At vanishing densities $n_{\mathrm{gas}}\rightarrow 0$, as indicated by $B_{0}$ in Table 3, in-medium clusters become more bound than those in vacuum due to the reduction of Coulomb energy Xia (2025). The inclusion of $\Lambda$ hyperons also increase the binding energy by $B_{\Lambda}$ and $B_{\Lambda\Lambda}$ for single and double $\Lambda$ hyperclusters. As $n_{\mathrm{gas}}$ increases, the binding energies of clusters decrease, which would eventually become unstable and melt at sufficient large densities, i.e., Mott transition. Despite the increased binding energy for $\Lambda$ hyperclusters at small $n_{\mathrm{gas}}$, the critical densities for vanishing binding energy $B=0$ are smaller for $\Lambda$ hyperclusters than that of normal clusters without $\Lambda$ hyperons. This is attributed to the reduction of binding energy per baryon for light nuclei with addition of $\Lambda$ hyperons, i.e., $B_{\mathrm{core}}/(A-1)>B_{\Lambda}$ as indicated in Fig. 1. Meanwhile, the variation of binding energies with respect to $Y_{p}$ are found to be sensitive to the differences between neutron and proton numbers in the cluster. For (hyper)clusters 3H, $^{4}_{\Lambda}$H, and $^{5}_{\Lambda\Lambda}$H, their binding energies decrease with the proton fraction of nuclear medium $Y_{p}$. If we replace a neutron in 3H by a $\Lambda$ hyperon, as indicated in Fig. 3, the binding energy for $^{3}_{\Lambda}$H is generally insensitive to $Y_{p}$ since $N_{p}=N_{n}=1$. Similarly, the binding energy for 4He, $^{5}_{\Lambda}$He, and $^{6}_{\Lambda\Lambda}$He also vary little with respect to the variation of $Y_{p}$ since $N_{p}=N_{n}=2$.
@@
$^{3}_{\Lambda}$H 5.208 -3.380 $195$ -0.0131 -0.31 $0.49$
3H 11.53 -3.089 $113$ -1.11 -1.2 $1.1$
$^{4}_{\Lambda}$H 13.48 -3.822 $144$ -1.19 -1.1 $1.2$
$^{5}_{\Lambda\Lambda}$H 15.31 -4.580 $170$ -1.19 -1.1 $1.2$
3He 10.57 -4.388 $149$ $1.197$ -0.26 $1.0$
$^{4}_{\Lambda}$He 12.70 -5.208 $182$ $1.377$ -0.34 $0.81$
4He 27.70 -5.029 $122$ $0.166$ -0.60 $0.77$
$^{5}_{\Lambda}$He 30.39 -5.800 $141$ $0.233$ -0.53 $0.51$
$^{6}_{\Lambda\Lambda}$He 32.97 -6.577 $162$ $0.285$ -0.47 $0.30$
5Li 26.12 -6.076 $219$ $0.374$ -0.11 $1.3$
$^{6}_{\Lambda}$Li 30.47 -7.383 $275$ $0.512$ 0.16 $1.6$
$^{7}_{\Lambda\Lambda}$Li 34.35 -8.371 $266$ $0.915$ 0.024 $1.6$
8Be 43.34 -10.06 $262$ $0.268$ $0.10$ 0
$^{9}_{\Lambda}$Be 50.80 -10.66 $230$ $0.438$ $0.049$ 0
12C 95.40 -14.55 $221$ $1.03$ -0.37 0
$^{13}_{\Lambda}$C 106.9 -15.44 $249$ $1.03$ -0.34 0
16O 134.6 -19.32 $337$ $1.41$ -0.44 0
$^{17}_{\Lambda}$O 145.6 -20.19 $337$ $1.47$ -0.47 0
40Ca 383.0 -48.47 $714$ $4.53$ -1.3 0
$^{41}_{\Lambda}$Ca 399.7 -49.27 $723$ $4.69$ -1.5 0
48Ca 454.2 -52.32 $712$ -4.81 -3.4 0
$^{49}_{\Lambda}$Ca 473.1 -53.08 $715$ -4.61 -3.5 0
@@
The best-fit coefficients are then indicated in Table 3. The coefficient $B_{0}$ represent the binding energy of (hyper)clusters in nuclear medium at vanishing $n_{\mathrm{gas}}$. The coefficients $a$ and $b$ indicate the density dependent behavior for the binding energy of clusters in nuclear medium. Despite the increased binding energy with the inclusion of $\Lambda$ hyperons, $a$ becomes more negative and leads to a faster decline for the binding energy as $n_{\mathrm{gas}}$ increases. The coefficients $c$, $d$, and $f$ are connected to the variation of binding energy with respect to the isospin asymmetry of nuclear medium, which are dominated by the differences between neutron and proton numbers in the cluster. For example, the binding energies for $^{3}_{\Lambda}$H, 4He, $^{5}_{\Lambda}$He, and $^{6}_{\Lambda\Lambda}$He ($N_{p}=N_{n}$) are roughly insensitive to $Y_{p}$. The binding energies for 3He, 5Li, $^{6}_{\Lambda}$Li, and $^{7}_{\Lambda\Lambda}$Li ($N_{p}>N_{n}$) increase with $Y_{p}$, while those of 3H, $^{4}_{\Lambda}$H, $^{5}_{\Lambda\Lambda}$H, 48Ca, and $^{49}_{\Lambda}$Ca ($N_{p}<N_{n}$) decrease with $Y_{p}$ as $c$ takes negative values. This is mainly attributed to the isovector effects in nuclear medium with the differences in the mean-field potentials of protons and neutrons change as we vary $Y_{p}$.
Figure 4: Binding energy per nucleon for various (hyper)clusters emersed in SNM and PNM as functions of the density of nuclear medium $n_{\mathrm{gas}}$, which are obtained by employing the relativistic density functional DD-LZ1 Wei et al. (2020). The thin curves indicate normal clusters and thick curves correspond to those with one additional $\Lambda$ hyperon.
In Fig. 4 we present the binding energy per baryon for heavier (hyper)clusters emersed in SNM and pure neutron matter (PNM, $Y_{p}=0$) as functions of $n_{\mathrm{gas}}$. Similar to light clusters, the binding energies of (hyper)clusters decrease with $n_{\mathrm{gas}}$. Nevertheless, since the binding energy per baryon increase for $\Lambda$ hyperclusters with $N_{p}\geq 4$, they become more bound with the Mott transition densities increased. By comparing the upper and lower panels in Fig. 4, the variations of binding energies with respect to $Y_{p}$ can be identified, where ($\Lambda$ hyper)clusters with $N_{p}=N_{n}\lesssim 6$ are generally insensitive to the variations in $Y_{p}$. By fitting the binding energies indicated in Fig. 4 with Eq. (36), we can fix the corresponding coefficients, which are indicated in Table 3. The variations of binding energies with respect to $n_{\mathrm{gas}}$ and $Y_{p}$ can then be reproduced.
@@
The charge radii $R_{\mathrm{ch}}$ for the (hyper)clusters as functions of the density of nuclear medium $n_{\mathrm{gas}}$ are presented in Fig. 5, which correspond to those in Fig. 4 and are obtained with the relativistic density functional DD-LZ1 Wei et al. (2020). It is found that $R_{\mathrm{ch}}$ are generally increasing with $n_{\mathrm{gas}}$ while the binding energy is decreasing. Similar to the cases in vacuum as indicated in Table 2, the addition of a $\Lambda$ hyperon reduces the charge radii. The slope of $R_{\mathrm{ch}}$ as a function of $n_{\mathrm{gas}}$ also becomes smaller with the addition of a $\Lambda$ hyperon. By comparing the upper and lower panels in Fig. 5, the isovector effects in nuclear medium can be identified, e.g., $R_{\mathrm{ch}}$ increases slightly faster for (hyper)clusters with in PNM than those in SNM as we increase $n_{\mathrm{gas}}$. For heavy (hyper)clusters such as 48Ca and $^{48}_{\Lambda}$Ca, the variations of $R_{\mathrm{ch}}$ with respect to $Y_{p}$ are less significant.
@@
Figure 6: Binding energies of 3H (thin curves), 4He (thin curves), $^{4}_{\Lambda}$H (thick curves) and $^{5}_{\Lambda}$He (thick curves) emersed in SNM and PNM as functions of $n_{\mathrm{gas}}$, which are obtained adopting the relativistic density functionals PK1 Long et al. (2004), TM1 Sugahara and Toki (1994b), and DD-LZ1 Wei et al. (2020). The results are fitted with Eq. (36) with the coefficients indicated in Table 4.
@@
Finally, to show the variations due to the uncertainties in relativistic density functionals, we examine the binding energies of 3H, 4He, $^{4}_{\Lambda}$H, and $^{5}_{\Lambda}$He emersed in SNM and PNM adopting the functionals PK1 Long et al. (2004), TM1 Sugahara and Toki (1994b), and DD-LZ1 Wei et al. (2020), which are presented in Fig. 6. As expected, the binding energy increases with the inclusion of $\Lambda$ hyperon, while the binding energy per baryon decreases since $B_{\mathrm{core}}/(A-1)>B_{\Lambda}$ for light hyperclusters as indicated in Fig. 1. The binding energies of the (hyper)clusters are generally decreasing with the density of nuclear matter $n_{\mathrm{gas}}$. For the cluster 4He, the binding energy decline slowly with $n_{\mathrm{gas}}$, while slight variations on the in-medium binding energy shift are observed when we adopt different functionals. Due to the non-linear self-couplings of mesons, the binding energies of 3H, $^{4}_{\Lambda}$H, and $^{5}_{\Lambda}$He decrease drastically at $n_{\mathrm{gas}}\approx 0.0005$-0.0011 $\mathrm{fm}^{-3}$ when we adopt non-linear relativistic density functionals PK1 Long et al. (2004) and TM1 Sugahara and Toki (1994b), while the density-dependent functional DD-LZ1 Wei et al. (2020) predicts relatively slow decline of the binding energies. It is worth mentioning that the binding energy of $^{5}_{\Lambda}$He decreases drastically when non-linear relativistic density functionals are adopted, even though that of 4He decreases slowly with $n_{\mathrm{gas}}$. In such cases, the addition of $\Lambda$ hyperons destabilize the light (hyper)clusters which is particular the case when non-linear relativistic density functionals are adopted.
@@
Table 4: Coefficients for the expansion formula (36) of binding energies for in-medium 3H, 4He, $^{4}_{\Lambda}$H, and $^{5}_{\Lambda}$He presented in Fig. 6, where $f=0$ is assumed.